# Transforming Strategic Games into Biform Games: Applications in Allocation Mechanisms and Green Technology Investment

### v1

@@ 2.4 Give examples to illustrate
3 Example Three monks have no water to drink (Chinese proverb)
The proverb goes “One monk will get two buckets of water, two monks will share a load of water, three monks will have no water.”
The proverb suggests that with many people, no one wants to take the major responsibility. It can be related to the American proverb: “Too many cooks in the kitchen.” With too many people, it’s hard to divide responsibility because no one would feel right with someone resting.
One monk lived on a mountain, and every day he went down the mountain with a stick with two buckets to haul water. However one day, another monk came to visit. These two monks carry two buckets of water together. When the third monk arrived, they had no incentive to carry water.
Let $\gamma>0$ represent the number of buckets of water and $w\left(\gamma\right)>0$ represent the utility derived by each monk from $\gamma$ buckets of water consumption, assuming $w\left(0\right)=0$ and $w\left(\gamma\right)$ an increasing function. Let $b\left({{\text{\;}}\gamma}\right),$ $c\left({{\text{\;}}\gamma}\right)>0$ denote the labor cost for one monk to retrieve $\gamma$ buckets of water and the labor cost for two monks share the task of fetching two buckets of water. Assume that the marginal utility of water is decreasing and $w\left(2\right)-b\left(2\right)>0,\;c\left(2\right)<b\left(2\right)$, $w\left(1\right)<b\left(2\right)$, $w\left(2\right)-w\left({\frac{2}{3}}\right)<\frac{1}{3}b\left(2\right)$.
Let $N$ denote the set monks, $i\in N$ denote Monk $i$ and $X_{i}$} denote the strategy set for Monk $i$. For each $i\in N$, ${X_{i}}=\{{s_{i1}},\;{s_{i2}}\}$ that represents each monk have two choses carrying water and inaction, where $s_{i1}$} denotes carrying water and $s_{i2}$} denotes inaction. If these two monks are uncooperative with each other, i.e., each monk is solely responsible for his water consumption. This story can be reprinted as a strategic game $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right)$.
A.The case of one monk.
If one monk lived on a mountain, he has two option of either inaction or carrying water. It is easy to check that the monk will chose carrying water individually as $w\left(2\right)-{\text{\;}}b\left(2\right)>0=w\left(0\right)$. This means “One monk will get two buckets of water.”
If $N=\left\{{1,2}\right\}$ or $N=\left\{{1,2,3}\right\}$, let $i\in N$ represent Monk $i$, $X_{i}$} denote the strategy set for Monki\;i. Here, $s_{i1}$} represents the action of carrying water, while $s_{i2}$} represents inaction. If the monks act independently, meaning each monk is solely responsible for their own water consumption, then in both scenarios, the behavior of each monk remains consistent, similar to the case with a single monk. We will analyze the two-monk scenario as an illustrative example. This situation can be modeled as a noncooperative game $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({{g_{i}}}\right)}\right)$ (Table LABEL:22).
Table 2: Two monks are uncooperative with each other
Monk 1
$S_{11}$ $S_{12}$
Monk 2 $S_{21}$ w(2)b(2),w(2)b(2)w(2)-b(2),\ w(2)-b(2) w(0),w(2)b(2)w(0),\ w(2)-b(2)
$S_{22}$ w(2)b(2),w(0)w(2)-b(2),\ w(0) w(0),w(0)w(0),\ w(0)
Note that $w\left(2\right)-b\left(2\right)>0=w\left(0\right)$. It is clear that $\left({{s_{11}},{s_{21}}}\right)$ is the unique Nash equilibrium. The situation can also be regarded as a direct extension of the individual monk scenario, wherein each monk is independently responsible for both the procurement and utilization of water.
B. The cases of two monks.
[…]
In the following section, we examine the impact resulting from the allocation of cooperation profits, with particular emphasis on the methodology for sharing labor costs. Assume that two monks share alike the water whoever has carry it, but they do not share the labor costs. Then, it is described as a strategic game $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({{f_{i}}}\right)}\right).$
$\begin{split}{f_{1}}\left({{s_{11}},{s_{21}}}\right)={f_{2}}\left({{s_{11}},{s_{21}}}\right)=w\left(2\right)-b\left(2\right);\\ {f_{1}}\left({{s_{11}},{s_{22}}}\right)=w\left(1\right)-b\left(2\right),{f_{2}}\left({{s_{11}},{s_{22}}}\right)=w\left(1\right);\\ {f_{1}}\left({{s_{12}},{s_{21}}}\right)=w\left(1\right),{f_{2}}\left({{s_{12}},{s_{21}}}\right)=w\left(1\right)-b\left(2\right);\\ {f_{1}}\left({{s_{12}},{s_{22}}}\right)={f_{2}}\left({{s_{12}},{s_{22}}}\right)=0.\end{split}$
The function ${v_{\rho}}:{2^{N}}\times X\to R\;$ is the characteristic function with strategy profiles, where $u_{f}(S,\cdot)=\textstyle\sum_{i\in S}f_{i}(\,\cdot\,)$.Then
$\begin{split}{v_{f}}\left({\emptyset,\left({{s_{ij}},{s_{ij}}}\right)}\right)&=\;{u_{f}}\left({\emptyset,\left({{s_{ij}},{s_{ij}}}\right)}\right)=0,i,j\in\{1,2\}\\ {v_{f}}\left({S,\left({{s_{11}},{s_{21}}}\right)}\right)&=w\left(2\right)-b\left(2\right)=\;{u_{f}}\left({S,\left({{s_{11}},{s_{21}}}\right)}\right),S=\left\{1\right\},\left\{2\right\};\\ {v_{f}}\left({\left\{{1,2}\right\},\left({{s_{11}},{s_{21}}}\right)}\right)&=w\left(1\right)+w\left(1\right)-c\left(2\right)>2\left({w\left(2\right)-b\left(2\right)}\right)={u_{f}}\left({\left\{{1,2}\right\},\left({{s_{11}},{s_{21}}}\right)}\right);\\ {v_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right)&=w\left(1\right)-b\left(2\right)=\;{u_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right);\\ {v_{f}}\left({\left\{2\right\},\left({{s_{11}},{s_{22}}}\right)}\right)&=w\left(1\right)=\;{u_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right);\\ {v_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right)&=w\left(1\right)-b\left(2\right)=\;{u_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right);\\ {v_{f}}\left({\left\{2\right\},\left({{s_{11}},{s_{22}}}\right)}\right)&=w\left(1\right)=\;{u_{f}}\left({\left\{1\right\},\left({{s_{11}},{s_{22}}}\right)}\right);\\ {v_{f}}\left({S,\left({{s_{12}},{s_{22}}}\right)}\right)&={u_{f}}\left({S,\left({{s_{12}},{s_{22}}}\right)}\right)=0,\;\;S=\left\{1\right\},\left\{2\right\},\left\{{1,2}\right\}.\end{split}$
The allocation function of $\left({\left\{{1,2}\right\},{v_{f}},X}\right)$ denote as ${a^{{v_{f}}}}\left(s\right)=\left({a_{1}^{{v_{f}}}\left(s\right),a_{2}^{{v_{f}}}\left(s\right)}\right)$ which is a solution of the cooperative stage $\left({\left\{{1,2}\right\},{v_{f}}\left({\cdot,x}\right)}\right)$. If the allocation $a^{{v_{f}}}$} of coalition benefit dose not considers sharing the labor cost except two monks carry water together, the stage of noncooperative game of biform game $\left({N,{v_{f}},X}\right)$ denoted by $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$ is described as Table LABEL:23.
Table 3: Two monks collaborate with each other without sharing labor costs
Monk 1
$s_{11}$ $s_{12}$
Monk 2 $s_{21}$ $\frac{1}{2}c(2),\ w(1)-\frac{1}{2}c(2) w(1),w(1)$b(2)w(1),\ w(1)-b(2)
$s_{22}$ w(1)b(2),$w(1)$-b(2),\ w(1) 0, 00,\ 0
Note that $w\left(1\right)-b\left(2\right)<0$. It implies that (${s_{12}},{s_{22}})$ is the unique Nash equilibrium of the noncooperative game $\left({\left\{{1,2}\right\},\left({{X_{i}}}\right),\left({a_{i}^{{v_{f}}}}\right)}\right)$. In this scenario, no player is willing to undertake the task of carrying water, which highlights that when responsibilities are ambiguously defined, individuals may not be motivated to contribute effort.
If the allocation of coalition benefit not only considers sharing the water but also cares sharing the labor cost, the stage of noncooperative game of biform game $\left({N,{v_{f}},X}\right)$ is described as Table LABEL:24.
Table 4: Two monks collaborate with each other and share the associated labor costs
Monk 1
$S_{11}$ $S_{12}$
Monk 2 $S_{21}$ $w(1)-\frac{1}{2}c(2)$, $w(1)-\frac{1}{2}c(2)$ $w(1)-\frac{1}{2}b(2)$ , $w(1)-\frac{1}{2}b(2)$
$S_{22}$ $w(1)-\frac{1}{2}b(2)$ , $w(1)-\frac{1}{2}b(2)$ $w(0),w(0)$
In this context, $a_{1}^{{v_{f}}}\left({{s_{11}},{s_{22}}}\right)=a_{2}^{{v_{f}}}\left({{s_{11}},{s_{22}}}\right)=a_{1}^{{v_{f}}}\left({{s_{12}},{s_{21}}}\right)=a_{2}^{{v_{f}}}\left({{s_{12}},{s_{21}}}\right)=w\left(1\right)-\frac{1}{2}b\left(2\right)$ indicates that the two monks have reached a mutual agreement, such as alternating the responsibility of carrying water on a daily basis or offering compensation to the monk who carry water. It is evident that (${s_{11}},{s_{21}}$) constitutes the unique Nash equilibrium, indicating that both monks opt for collaboratively fetching one bucket of water. This outcome exemplifies the mutually beneficial results of cooperation. As the proverb says: “Two monks will share a load of water. ”
B.The cases of three monks.
Let $N=\{1,2,3\}$ represent the set of monks, where $i\in N$ denotes Monk $i$. ${S_{i}}=\left\{{{s_{i1}},{s_{i2}}}\right\}$ also represents the strategy set for Monk $i$, with $s_{i1}$} indicating the action of carrying water and $s_{i2}$} representing inaction.
[…]

### v2

@@ 2.4 Give examples to illustrate
3 Biform Game Analysis of Incentives for Multi-Departmental Regulatory Cooperation
3.1 Problem Statement
Food safety is of vital importance to the national economy and people¡¯s livelihoods. Strengthening food safety supervision and improving regulatory efficiency are the responsibilities of the government and an important part of the modernization of the national governance system. Since the advent of the new era, the Party and government have attached great importance to food safety supervision. However, food safety incidents still occur from time to time. Looking into the root causes, the problems are complex and the reasons are diverse, but the lack of effective cooperative supervision is also one of the important reasons. China has long implemented a segmented regulatory system, with highly dispersed regulatory powers, which has led to regulatory fragmentation. Regulatory performance depends not only on the performance of individual departments but also on their ability to cooperate. Even after the institutional reform of the merger of three bureaus in 2018, many regulatory matters still rely on the cooperation between the State Administration for Market Regulation and other ministries and bureaus, and also depend on the cooperation between subordinate departments. From the local perspective, in addition to cooperation between departments, cross-regional cooperation between local governments is also indispensable.
Food safety issues are characterized by their distinct cross-boundary nature. This is reflected not only in the boundaries of knowledge and jurisdiction, but also more prominently in the complex and diverse range of stakeholders. The political interests of regulatory agencies, the economic interests of the regulated entities, and the social interests of the public are interwoven, creating significant tensions. There is a need to promote collaboration among multiple stakeholders, to build mutual trust in the process, and thereby to integrate their interests to improve regulatory effectiveness. Therefore, strengthening cooperative regulation is seen as an important measure to resolve the regulatory dilemma.
Under this circumstance, collaborative supervision in the field of food safety has become a research focus for academia, and scholars have made extensive explorations in promoting collaborative supervision. Many scholars have drawn nourishment from the perspective of holistic government and delved into the specific field of food safety supervision. Xu [28] proposed that in China, the forms of collaborative supervision of food safety include cooperation among government agencies, between government and non-governmental organizations, and among non-governmental organizations. However, cooperation among government agencies is the most favored, and unitary government supervision still dominates. Ma [29] argued that, considering the costs of integration, national conditions, and the complexity of food safety issues, a single-agency system may not be able to effectively solve all the problems of China’s multi-departmental supervision. Yan and Nie [30] took relational contracts as the entry point and used the theory of transaction costs to explore the mechanism of the cooperation dilemma among China’s food safety supervision departments. Yan [31] suggested that improvements in departmental cooperation should be made in terms of organizational structure, responsibility and incentive mechanisms, partnerships, and organizational culture.
From the perspective of institutional economics, food safety supervision is a public good, and government supervision essentially represents the exercise of public property rights. The government and relevant departments perform their supervisory duties as a means of using public power to maintain food safety and to intervene in and constrain the food market (see Chen [32], Chen [33], and Chen [34]). Public power originates from the delegation and authorization by the public. Given that public property rights are characterized by indivisibility, non-exclusivity of usage rights, and externality, the exercise of supervisory power over them is also subject to the ”free-rider” phenomenon. Therefore, the public is reluctant to exercise this power individually and instead places it in the public domain, making it public power. Consequently, it is inevitable that the state or government acts on behalf of the public to exercise the public property right of food safety supervision.
The research results of Parisi et al. [35] indicate that when public property rights are jointly owned by multiple departments with equal power status and no hierarchical relationships, the efforts each department makes to maximize its own interests are greater than those required to maximize collective interests. This insufficient cooperation among departments leads to an inadequate supply of public goods, resulting in the Tragedy of the Anticommons.The concept of the Tragedy of the Anticommons was first proposed by Heller [36] from the University of Michigan. Anticommons refers to a situation where a resource or property has many owners, each of whom has the formal or informal right to prevent others from using the resource, ultimately leading to no one having effective and substantive usage rights. The Tragedy of the Anticommons is in contrast to the Tragedy of the Commons,which describes the dilemma caused by the overconsumption of public resources. The Tragedy of the Anticommons reveals that the underutilization of public resources (power) can also lead to idleness and waste.
In the field of food quality and safety supervision, the Tragedy of the Anticommons is the root cause of the low efficiency of collaborative supervision. This is mainly due to the government granting public resources (power) to different authorities. Numerous regulatory departments, burdened with overlapping responsibilities, fail to perform their duties effectively and even shirk responsibility. In China, food quality and safety regulatory power was once divided among nearly 10 departments, including industry and commerce, health, drug supervision, quality supervision, agriculture, commerce, and customs. These departments were responsible for food raw materials, processing, production, distribution, and consumption, and each was granted certain regulatory powers. However, no single department had comprehensive regulatory responsibilities. With multiple authorities involved, each acting independently and even setting obstacles to pursue its own interests, collaboration became extremely difficult. Some regulatory responsibilities were lost in the process of mutual shirking, leading to the idleness and waste of public power and even regulatory failure. In March 2018, the Central Committee of the Communist Party of China issued the ”Plan for Deepening the Reform of Party and State Institutions.” This plan integrated the responsibilities of the State Administration for Industry and Commerce, the General Administration of Quality Supervision, Inspection and Quarantine, the China Food and Drug Administration, the price supervision and anti-monopoly enforcement responsibilities of the National Development and Reform Commission, the anti-monopoly enforcement of business concentration of the Ministry of Commerce, and the Office of the State Council Anti-monopoly Commission to form the State Administration for Market Regulation. The State Administration for Market Regulation is responsible for the management and supervision of food safety. Its main responsibilities include formulating and implementing food safety policies, supervising food production and business activities, and issuing food safety standards and regulations to ensure food safety.
Does this mean that the much-criticized ”nine dragons managing water” situation in the field of food safety has come to an end? The reality is not so simple. According to the ”Food Safety Law of the People’s Republic of China” (hereinafter referred to as the ”Food Safety Law”) (revised for the second time by the Standing Committee of the 13th National People’s Congress on April 29, 2021) and other relevant laws, regulations, rules, and normative documents, in addition to the food department, departments such as agriculture, health, grain, environmental protection, public security, and industry and information technology still bear relevant functions of food safety supervision. That is to say, after March 2018, although the State Administration for Market Regulation is responsible for the supervision and management of food production and business activities, the multi-departmental supervision structure in China’s food safety field has not been completely changed. Many areas still require coordinated cooperation and joint management by multiple departments. To strengthen the coordination and cooperation of food safety supervision departments, the State Council Food Safety Commission is responsible for coordinating the food safety supervision work of various departments. However, how to strengthen cooperative supervision and how to eliminate shirking and wrangling remain major challenges that food safety supervision must face.
According to the provisions of laws, regulations, and rules such as the Food Safety Law, as well as the specific practices of regulatory authorities, the methods of coordination and cooperation among regulatory authorities are diverse. These include seeking opinions (consultation), cooperation agreements (memoranda of understanding), joint decision-making (the Food Safety Law often uses expressions such as ¡°in conjunction with…¡± or ¡°organize and carry out…¡±), information notification, joint law enforcement, administrative assistance, case referral, joint meetings, and liaison officer meetings. To a certain extent, these activities have promoted communication and cooperation among regulatory authorities. However, they also suffer from inefficiencies and even ineffectiveness to varying degrees.
The lack of incentive mechanisms for cooperation is a significant flaw in collaborative supervision. In practice, some regulatory authorities, in order to better fulfill their duties, have voluntarily taken some coordination and cooperation actions, such as signing cooperation agreements or memoranda of understanding, and establishing joint meeting or liaison officer meeting systems. However, such voluntary cooperation often lacks legal binding force. Cooperation agreements (memoranda) may be terminated at any time due to one party¡¯s refusal to cooperate, and joint meetings may not be held regularly due to the non-cooperation of any party. In other words, whether the purpose of voluntary cooperation can be achieved largely depends on whether the cooperating parties are honest, trustworthy, and committed to fulfilling their obligations. This, in turn, is closely related to whether the regulatory authorities can benefit from cooperation and the extent of such benefits. Under the existing system, there is an imbalance in power and resource allocation among food safety regulatory authorities. Some departments always give more and gain less in the cooperation process (such as information notification), while some departments may be pure beneficiaries. In the absence of specific incentive mechanisms, the enthusiasm of the former for voluntary cooperation will inevitably be affected. Since 2018, the supervision and management of food production and business operations have been the responsibility of the Market Supervision and Administration Bureau, which has led to centralized and unified management of regulatory work. However, local Market Supervision and Administration Bureaus, which operate under the management of local governments, need to strengthen cooperative supervision and overcome the constraints of local protectionism. Coordination and cooperation among regulatory authorities cannot merely remain at the level of principled provisions and slogan-like advocacy; they require scientific and meticulous institutional design. In addition to institutional constraints, incentive mechanisms are equally important for stimulating the proactivity and enthusiasm of regulatory authorities in coordination and cooperation. Drawing from international experience, an effective coordination and cooperation mechanism must have clear cooperation goals and a performance evaluation system. In the performance evaluation system, not only should there be assessments of the regulatory effectiveness of individual departments, but evaluations of inter-departmental coordination and cooperation are also essential. Moreover, the weight assigned to these evaluations is crucial for promoting multi-departmental coordination and cooperation. Establishing a scientific and rational performance evaluation system is of great significance for improving regulatory efficiency. On the one hand, performance assessments ensure that regulatory authorities fulfill their departmental responsibilities, which is consistent with the current evaluation mechanisms. On the other hand, by making inter-departmental coordination and cooperation an important part of the evaluation, regulatory authorities are encouraged to actively participate in collaborative supervision. Only in this way can a positive situation be created where departments strive to maximize their own interests by focusing on their main tasks while actively coordinating and cooperating in joint supervision. Therefore, the coexistence of competition and cooperation is the norm in multi-departmental supervision.
This section applies the biform game model established in Section 2 to the cooperative supervision of food quality and safety. It examines the gains and costs associated with departments’ inputs into cooperative supervision, as well as the distribution of cooperative gains, to investigate the strategic choices of participating departments. Specifically, it conducts an in-depth analysis of the competition and cooperation among regulatory authorities through the lens of the biform game. Focusing on the example of cooperative supervision involving three departments, it constructs a basic strategic game model. The strategic game is then treated as a two-stage biform game. By selecting appropriate cooperative gain distribution functions and analyzing the corresponding solutions of the biform game, it explores the strategic behaviors of regulatory authorities. Additionally, it offers suggestions and insights for the development of incentive mechanisms and the promotion of cooperative supervision.
3.2 Problem Description
In one area of food safety supervision, three supervisory departments need to cooperate. These departments may be agencies or departments within a single region involved in this area of supervision, or they may be cross-regional agencies or departments. Denote the set of supervisory departments as $N=\big\{1,2,3\big\}$. For each supervisory department $i\in N$, there are two strategic choices: to participate in collaborative supervision or not to participate in collaborative supervision, denoted as ”pp” and ”npnp”, respectively. The set of strategies for a supervisory department is represented by $S_{i}=\big\{p,np\big\}$. For the collaborative supervision task, the higher-level department or local government provides corresponding inputs or performance recognition. These inputs and performance are regarded as the utilities of the supervisory departments.
If local governments or higher-level authorities only conduct an overall assessment of the work without evaluating individual departments, the total utility cannot be disaggregated among the participating departments; instead, it must be split equally among the three. Under this rule, should a single department take on and complete the entire cooperative-supervision task, it would receive only one-third of the total benefit. Because the cost of acting alone exceeds this share, the department¡¯s net utility would be negative. If two departments work together, each still receives one-third of the total benefit; although collaboration lowers the aggregate cost, the benefit accruing to each participant remains below its individual cost, so their utilities are still negative. Only when all three departments participate does each again receive one-third of the total, but the cost falls further, so that each department¡¯s benefit exceeds its individual cost and its net utility becomes positive.
To construct the game model, we make the following assumptions:
A1. The payoff for completing the entire collaborative supervision task is $R$.
A2. The cost for a single department to complete the collaborative supervision task alone is $C$, satisfying $R>C$ and $\frac{R}{2}<C$. In the absence of cooperation between departments, if two departments participate and complete the collaborative supervision task, the cost $C$ is equally shared between the two departments. If three departments participate and complete the task, the cost $C$ is equally shared among the three departments.
A3. Considering the efficiency improvement due to the synergistic effect, if two departments collaborate to complete the collaborative supervision task, the combined cost for the two departments is $rC<C$; if three departments collaborate to complete the task, the combined cost for the three departments is $qC<rC<C$.
A4. Regarding the collaborative supervision task, the utility of a supervisory department is the payoff minus the cost.
To maximize their own utility, there is competition among the departments. Meanwhile, in regulatory fields involving multiple departments, cooperation is needed to improve efficiency. Therefore, competition and cooperation coexist. The above problem can be analyzed using the dual-type game discussed in Section 2.
3.3 Non-Cooperative Scenarios in Strategy Games
Let $S=\prod_{i\in N}S_{i}$} be the set of strategy profiles. For each $i\in N$, the utility of the supervisory department is as follows:
$\begin{split}\begin{aligned} &\pi_{{i}}({np,np,np})=0,\\ &\pi_{1}({p,np,np})=\frac{1}{3}{R-C,}\ \pi_{2}({p,np,np})=\frac{1}{3}{R,}\ \pi_{3}({p,np,np})=\frac{1}{3}{R,}\\ &\pi_{1}({np,p,np})=\frac{1}{3}{R,}\ \pi_{2}({np,p,np})=\frac{1}{3}{R-C,}\ \pi_{3}({np,p,np})=\frac{1}{3}{R,}\\ &\pi_{1}({np,np,p})=\frac{1}{3}{R,}\ \pi_{2}({np,np,p})=\frac{1}{3}{R,}\ \pi_{3}({np,np,p})=\frac{1}{3}{R-C,}\\ &\pi_{1}({np,p,p})=\frac{1}{3}{R,}\ \pi_{2}({np,p,p})=\frac{1}{3}{R-\frac{1}{2}C,}\ \pi_{3}({np,p,p})=\frac{1}{3}{R-\frac{1}{2}C,}\\ &\pi_{1}({p},{n}{p},{p})=\frac{1}{3}{R}-\frac{1}{2}{C},\ \pi_{2}({p},{n}{p},{p})=\frac{1}{3}{R},\ \pi_{3}({p},{n}{p},{p})=\frac{1}{3}{R}-\frac{1}{2}{C},\\ &\pi_{1}({p},{p},{n}{p})=\frac{1}{3}{R}-\frac{1}{2}{C},\ \pi_{2}({p},{p},{n}{p})=\frac{1}{3}{R}-\frac{1}{2}{C},\ \pi_{3}({p},{p},{n}{p})=\frac{1}{3}{R},\\ &\pi_{i}({p,p,p})=\frac{1}{3}{R-\frac{1}{3}C.}\end{aligned}\end{split}$
Therefore, the cooperative regulation problem can be described as a strategic game with $(N,(S_{i}),(\pi_{i})),\ \pi_{i}:S\to R$} as the payoff functions of the supervisory department $i$.
To take into account the participation level or the probability of participation of the supervisory departments, we directly discuss the mixed game over $(N,(S_{i}),(\pi_{i}))$. For any $i\in N$, let
$\Delta_{i}=\big\{u_{i}=(x_{i};1-x_{i});\ x_{i}\in[0,1]\big\}$
and $\Delta=\prod_{i\in N}\Delta_{i}$}. Then $\Delta_{i}$ represents the set of mixed strategies for $i$, and $\Delta$ represents the set of mixed strategy profiles for the game. Here, $x_{i}$ denotes the degree or probability of department participating in cooperative regulation, and correspondingly, $1-x_{i}$ denotes the degree or probability of non-participation. The payoff function for the mixed game is
$\begin{aligned} f_{1}(u_{1},u_{2},u_{3})&=x_{1}x_{2}x_{3}\pi_{1}(p,p,p)+x_{1}x_{2}(1-x_{3})\pi_{1}(p,p,np)+x_{1}(1-x_{2})x_{3}\pi_{1}(p,np,p)\\ &+(1-x_{1})x_{2}x_{3}\pi_{1}(np,p,p)+x_{1}(1-x_{2})(1-x_{3})\pi_{1}(p,np,np)\\ &+(1-x_{1})x_{2}(1-x_{3})\pi_{1}(np,p,np)+(1-x_{1})(1-x_{2})x_{3}\pi_{1}(np,np,p)\\ &+(1-x_{1})(1-x_{2})(1-x_{3})\pi_{1}(np,np,np)\\ &=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{3}\\ &-\frac{1}{3}Rx_{2}x_{3}+\left(\frac{1}{3}R-C\right)x_{1}+\frac{1}{3}R(x_{2}+x_{3}).\end{aligned}$
Similarly, for any $(u_{1},u_{2},u_{3})\in\Delta$, we can obtain
$\begin{aligned} f_{2}(u_{1},u_{2},u_{3})&=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{2}x_{3}\\ &-\frac{1}{3}Rx_{1}x_{3}+\frac{1}{3}Rx_{1}+\left(\frac{1}{3}R-C\right)x_{2}+\frac{1}{3}Rx_{3},\\ f_{3}(u_{1},u_{2},u_{3})&=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\frac{1}{3}Rx_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{2}x_{3}\\ &-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{3}+\frac{1}{3}R(x_{1}+x_{2})+\left(\frac{1}{3}R-C\right)x_{3}.\end{aligned}$
Since $u_{i}=(x_{i},1-x_{i})$, then for any $(x_{i},1-x_{i})\in\Delta_{i}$, the aforementioned payoff function can be expressed as
$\begin{aligned} f_{1}(x_{1},x_{2},x_{3})&=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{3}\\ &-\frac{1}{3}Rx_{2}x_{3}+\left(\frac{1}{3}R-C\right)x_{1}+\frac{1}{3}R(x_{2}+x_{3}),\\ f_{2}(x_{1},x_{2},x_{3})&=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{2}x_{3}\\ &-\frac{1}{3}Rx_{1}x_{3}+\frac{1}{3}Rx_{1}+\left(\frac{1}{3}R-C\right)x_{2}+\frac{1}{3}Rx_{3},\\ f_{3}(x_{1},x_{2},x_{3})&=\left(\frac{1}{3}R-\frac{1}{3}C\right)x_{1}x_{2}x_{3}-\frac{1}{3}Rx_{1}x_{2}-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{2}x_{3}\\ &-\left(\frac{1}{3}R-\frac{1}{2}C\right)x_{1}x_{3}+\frac{1}{3}R(x_{1}+x_{2})+\left(\frac{1}{3}R-C\right)x_{3}.\end{aligned}$ (7)
Eq. (7) constitutes a strategic game, denoted by $(N,(\Delta_{i}),(f_{i}))$. If we regard Eq. (7) as a non-cooperative game, we discuss the Nash equilibrium of this game. By taking the partial derivatives of $f_{1},f_{2},f_{3}$ with respect to $x_{1},x_{2},x_{3}$, respectively, we can obtain
$\begin{gathered}\frac{\partial{f}_{1}}{\partial{x}_{1}}=\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x}_{2}{x}_{3}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{2}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{3}+\left(\frac{1}{3}{R}-{C}\right),\\ \frac{\partial{f}_{2}}{\partial{x}_{2}}=\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x}_{1}{x}_{3}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{1}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{3}+\left(\frac{1}{3}{R}-{C}\right),\\ \frac{\partial{f}_{3}}{\partial{x}_{3}}=\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x}_{1}{x}_{2}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{2}-\left(\frac{1}{3}{R}-\frac{1}{2}{C}\right){x}_{1}+\left(\frac{1}{3}{R}-{C}\right).\end{gathered}$ (11)
It is noted that
$\begin{aligned} \frac{\partial^{\prime}{f_{1}}}{\partial{~x_{1}}}&=\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x_{2}}{x_{3}}-\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x_{2}}-\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right){x_{3}}+\frac{1}{6}{C}{x_{2}}+\frac{1}{6}{C}{x_{3}}+\left(\frac{1}{3}{R}-{C}\right)\\ &\leq-\left(\frac{1}{3}{R}-\frac{1}{3}{C}\right)[{x}_{2}+(1-{x}_{2}){x}_{3}]+\frac{1}{6}{C}+\left(\frac{1}{3}{R}-{C}\right)\\ &\leq\frac{1}{6}{C}+\left(\frac{1}{3}{R}-{C}\right)=\frac{1}{3}{R}-\frac{5}{6}{C}<\frac{1}{3}{R}-\frac{5}{12}{C}<0.\end{aligned}$
Similarly, we can obtain $\frac{\partial^{\prime}{f_{2}}}{\partial{~x_{2}}}<0$ and $\frac{\partial^{\prime}{f_{3}}}{\partial{~x_{3}}}<0$. From this, it can be seen that $f_{1}(x_{1},x_{2},x_{3}),f_{2}(x_{1},x_{2},x_{3}),f_{3}(x_{1},x_{2},x_{3})$} are decreasing functions with respect to $x_{1},x_{2},x_{3}$, respectively. Therefore, $(0,0,0)$ constitute the unique Nash equilibrium of the game (2.2.1). In this case, each department, aiming to maximize its own utility, will inevitably choose not to participate in cooperative regulation. This conclusion reflects the ¡°tragedy of the anti-commons¡± in food safety cooperative regulation, that is to say, when performing the function of cooperative regulation, it often falls into such a deadlock or dilemma, namely, the regulatory departments will often choose inaction in cooperative regulation.
It is worth noting that in the model given by Eq. (7), the total utility obtained by the three departments through joint cooperative regulation is higher than the total utility obtained by inaction, that is
$f_{1}(1,1,1)+f_{2}(1,1,1)+f_{3}(1,1,1)=\frac{1}{3}(R-C)>0=f_{1}(0,0,0)+f_{2}(0,0,0)+f_{3}(0,0,0).$
Furthermore, denote
${g(x_{1},x_{2},x_{3})=f_{1}(x_{1},x_{2},x_{3})+f_{2}(x_{1},x_{2},x_{3})+f_{3}(x_{1},x_{2},x_{3}).}$
Then,
[…]